Question 7 (Marks: 10) Using Nodal analysis, calculate the branch current ix.

5 Ω ΛΜ ΛΑ ΛΜ 3Ω 6 Ω (1) 4A 3ία 7 Ω.
Answer #1

lav,- 14V2 - 72t 73-642=0 352 72 31 ist At node y 4=ixtim » 3=V1 V2 TV1_V3 Multiplying by is 45=5V-5V2+3V ~ 3V3 su, -582-3v3=3 3ov, - 30v2-64 + 6V3-5V₂ +5V₂ = 0 ZAW-35V2 + 113=0 > solving equations , & i vi=2.4375, V2=-0.656250, N3=-7.40695 in=2.4375